A certain recipe asks for 7/9 cup of granola. How much is needed to make 1/9 of a recipe.
Answer by TimothyLamb(4379)

g = 7/9 * 1/9
g = (7*1)/(9*9)
g = 7/81
g ~= 0.0864 cups
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php